# Metaverse for Safer Roadways: An Immersive Digital Twin Framework for Exploring Human-Autonomy Coexistence in Urban Transportation Systems

### v1

@@ III-C Results and Discussion
In terms of interaction interfaces, user ratings again suggest a gradual increase in involvement across the 4 modalities. User adaptation/immersion for the first 3 configurations were rated almost similarly with a slightly higher rating for the fourth one. Finally, the interface quality ratings suggest the highest discomfort/interference for the first two configurations, followed by the third one, and least for the fourth one. It is worth mentioning that some users experienced occasional discomfort in locating certain interaction modalities (e.g. keys, cursor position, pedals) when wearing the HMD.
@@ IV-C Results and Discussion
Although all the configurations safely avoided collision with the virtual peer vehicle, their performance varied. The AV performed the most aggressive acceleration and braking but traveled the farthest before coming to a safe stop. This is likely due to the peer vehicle being detected only after it was captured in the perception field of view. The CAV, on the other hand, was able to detect the presence of the peer vehicle ahead of time through V2V communication and therefore able to stop much earlier and smoothly. In terms of manual driving, it was observed that the performance was greatly influenced by the observation interface. The user with static HMD configuration performed worse than the CAV due to the same issue of not being able to observe the peer vehicle in time. Conversely, the user with dynamic HMD configuration could observe the peer vehicle approaching by turning their head and therefore performed the best out of all the configurations in terms of stopping distance.

### v2

@@ III-C Results and Discussion
In terms of interaction interfaces, user ratings again suggest a gradual increase in involvement across the 4 modalities. User adaptation/immersion for the first 3 configurations were rated almost similarly with a slightly higher rating for the fourth one. Finally, the interface quality ratings suggest the highest discomfort/interference for the first two configurations, followed by the third one, and the least for the fourth one. It is worth mentioning that some users experienced occasional discomfort in locating certain interaction modalities (e.g. keys, cursor position, pedals) when wearing the HMD.
@@ IV-C Results and Discussion
Although all the configurations safely avoided collision with the virtual peer vehicle, their performance varied. The AV performed the most aggressive acceleration (2.35 m/s2) and braking (-7.59 m/s2) but traveled the farthest (110.21 m) before coming to a safe stop. This is likely because the peer vehicle was detected only after it was captured in the perception field of view. The CAV, on the other hand, was able to detect the presence of the peer vehicle ahead of time through V2V communication and therefore able to stop much earlier (105.49 m) and smoothly (-1.26 m/s2). In terms of manual driving, it was noted that the observation interface greatly influenced the performance. The user with static HMD configuration performed worse than the CAV due to the same issue of being unable to observe the peer vehicle in time. Conversely, the user with dynamic HMD configuration could observe the peer vehicle approaching by turning their head and therefore performed the best out of all configurations in terms of stopping distance (103.86 m).